find three consecutive odd integers such that the sum of the first, two times the second, and three times the third is 82
Found 2 solutions by checkley77, RAY100:
Answer by checkley77(12569)

Let x, x+2 & x+4 be the three integers.
x+2(x+2)+3(x+4)=82
x+2x+4+3x+12=82
6x=82-4-12
6x=66
x=66/6
x=11 the smallest integer.
11=2=13 for the next integer.
11=4=15 for the largest integer.
Proof:
11+2*13+3*15=82
11+26+45=82
82=82

Answer by RAY100(1637)

Let x=1st integer, x+2 = 2nd integer, and x+4 = 3rd integer
.
1(x) +2(x+2) +3(x+4) = 82
.
x+2x+4+3x+12 =82
.
6x + 16 =82
.
6x=66
.
x=11
.
Integers are 11,13,15,
.
.
check,,11 +2*13 +3*15 = 82,,,,,ok
.